# Examining Kempe equivalence via commutative algebra

### v1

@@ 5. Examining Kempe equivalence
$f_{1}(x)=\begin{cases}f(x)&x\notin H,\\ \mu&x\in H\mbox{ and }f(x)=\eta,\\ \eta&x\in H\mbox{ and }f(x)=\mu.\\ \end{cases}$
@@ 6. Computing Kempe classes via Kempe ideals
The set of all monomials in $M_{G}$ consists of all monomials associated to $k$-colorings of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{d}$ such that $G_{{\mathbf{a}}}$ is not an induced subgraph of $G$.
@@
Let $G$ be a graph on $[d]$, and let $\mathcal{G}_{1}$ be the reduced Gröbner basis of $J_{G}$ with respect to a monomial order $<$ and set $\mathcal{G}_{2}=\{x_{S}x_{T}\mid S,T\in S(G),\ S\cap T\neq\emptyset\}$. Then
@@
is the reduced Gröbner basis of $K_{G}$ with respect to $<$.
@@
Since $\mathcal{G}_{1}$ is a Gröbner basis of $J_{G}$, we have $J_{G}=\langle\mathcal{G}_{1}\rangle$. It follows from $M_{G}=\langle\mathcal{G}_{2}\rangle$ that ${\mathcal{G}}$ is a set of generators of $K_{G}=J_{G}+M_{G}$. We apply Buchberger Criterion to ${\mathcal{G}}$. If $p,q\in\mathcal{G}_{1}\setminus M_{G}$, then the remainder of $S(p,q)$ on division by ${\mathcal{G}}$ is 00 since $\mathcal{G}_{1}$ is a Gröbner basis of $J_{G}$. If $p,q\in\mathcal{G}_{2}$, then both $p$ and $q$ are monomials, and hence $S(p,q)=0$. Suppose that $p\in\mathcal{G}_{2}$ and $q\in\mathcal{G}_{1}\setminus M_{G}$. Let $p=x_{S}x_{T}$ with $S,T\in S(G)$ and $S\cap T\neq\emptyset$, and let $q={\mathbf{x}}_{f}-{\mathbf{x}}_{g}$ where $f$ and $g$ are $k$-colorings of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in{\mathbb{Z}}^{d}_{\geq 0}$ and ${\rm in}_{<}(q)={\mathbf{x}}_{f}$.
@@
Case 2. (${\mathbf{x}}_{f}$ is divided by $x_{S}$.) From Lemma 6.3, ${\mathbf{x}}_{f}$ is not divided by $x_{T}$ since $q\notin M_{G}$. Then $S(p,q)=x_{T}{\mathbf{x}}_{g}$. Let $i\in S\cap T$. Since ${\mathbf{x}}_{f}$ is divided by $x_{S}$, $\pi({\mathbf{x}}_{f})=\pi({\mathbf{x}}_{g})$ is divided by $t_{i}$. Thus $\pi(x_{T}{\mathbf{x}}_{g})$ is divided by $t_{i}^{2}$. Hence $x_{T}{\mathbf{x}}_{g}$ is divided by a monomial $x_{S^{\prime}}x_{T}$ such that $i\in S^{\prime}\cap T$. Then the remainder of $S(p,q)=x_{T}{\mathbf{x}}_{g}$ on division by ${\mathcal{G}}_{2}$ is 00.
Therefore the remainder of any $S$-polynomial on division by ${\mathcal{G}}$ is 00. By Buchberger Criterion, ${\mathcal{G}}$ is a Gröbner basis of $K_{G}$. Since ${\mathcal{G}}_{1}$ is reduced, it is easy to see that ${\mathcal{G}}$ is reduced. ∎
@@
Suppose that ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in K_{G}$. We show that ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in J_{G}$. Let $\mathcal{G}$ be the reduced Gröbner basis of $K_{G}$ with respect to a monomial order $<$. From Proposition 6.4, $\mathcal{G}=(\mathcal{G}_{1}\setminus M_{G})\cup\mathcal{G}_{2}$ where $\mathcal{G}_{1}$ is the reduced Gröbner basis of $J_{G}$ with respect to $<$ and $\mathcal{G}_{2}=\{x_{S}x_{T}\mid S,T\in S(G),\ S\cap T\neq\emptyset\}$. Let ${\mathbf{x}}_{f^{\prime}}$ (resp. ${\mathbf{x}}_{g^{\prime}}$) denote the remainder of ${\mathbf{x}}_{f}$ (resp. ${\mathbf{x}}_{g}$) on division by $\mathcal{G}_{1}\setminus M_{G}$. Suppose that ${\mathbf{x}}_{f^{\prime}}\neq{\mathbf{x}}_{g^{\prime}}$. Since $f$ and $g$ are $k$-colorings of $G_{0}$, we have ${\mathbf{x}}_{f},{\mathbf{x}}_{g}\notin M_{G}$ from Lemma 6.1. From Lemma 6.3, ${\mathbf{x}}_{f^{\prime}},{\mathbf{x}}_{g^{\prime}}\notin M_{G}$ since both ${\mathbf{x}}_{f}-{\mathbf{x}}_{f^{\prime}}$ and ${\mathbf{x}}_{g}-{\mathbf{x}}_{g^{\prime}}$ belong to $J_{G}$. Thus ${\mathbf{x}}_{f^{\prime}}-{\mathbf{x}}_{g^{\prime}}$ ($\neq 0$) is the normal form of ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in K_{G}$ with respect to $\mathcal{G}$. This contradicts that $\mathcal{G}$ is a Gröbner basis of $K_{G}$. Thus ${\mathbf{x}}_{f^{\prime}}={\mathbf{x}}_{g^{\prime}}$. Then the normal form of ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}$ with respect to $\mathcal{G}_{1}\setminus M_{G}$ is zero, and hence ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in J_{G}$. ∎
@@
Let $G$ be a graph on $[d]$ and $<$ a monomial order on $R[G]$, and let $\{{\mathbf{x}}_{f_{1}},\dots,{\mathbf{x}}_{f_{s}}\}$ be the set of all standard monomials of degree $k$ with respect to ${\rm in}_{<}(K_{G})$. Then each $f_{i}$ is a $k$-coloring of an induced subgraph of $G$. In addition, given an induced subgraph $G^{\prime}$ of $G$,
@@
Let $\mathcal{G}$ be the reduced Gröbner basis of $K_{G}$ with respect to a monomial order $<$. From Proposition 6.4, $\mathcal{G}=(\mathcal{G}_{1}\setminus M_{G})\cup\mathcal{G}_{2}$ where $\mathcal{G}_{1}$ is the reduced Gröbner basis of $J_{G}$ with respect to $<$ and $\mathcal{G}_{2}=\{x_{S}x_{T}\mid S,T\in S(G),\ S\cap T\neq\emptyset\}$. Since each ${\mathbf{x}}_{f_{i}}$ is not divided by any monomial in $\mathcal{G}_{2}$, from Lemma 6.1, each $f_{i}$ is a $k$-coloring of an induced subgraph of $G$.
@@
Let $f$ be a $k$-coloring of $G^{\prime}$, and let ${\mathbf{x}}_{f^{\prime}}$ be the remainder of ${\mathbf{x}}_{f}$ on division by $\mathcal{G}_{1}\setminus M_{G}$. Then ${\mathbf{x}}_{f}-{\mathbf{x}}_{f^{\prime}}$ belongs to $J_{G}$. Since ${\mathbf{x}}_{f}\notin M_{G}$, we have ${\mathbf{x}}_{f^{\prime}}\notin M_{G}$ from Lemma 6.3. Thus ${\mathbf{x}}_{f^{\prime}}$ is the normal form of ${\mathbf{x}}_{f}$ with respect to $\mathcal{G}$, and hence ${\mathbf{x}}_{f^{\prime}}={\mathbf{x}}_{f_{j}}$ for some $j$. Since ${\mathbf{x}}_{f}-{\mathbf{x}}_{f_{j}}$ belongs to $J_{G}$, we have $f\sim_{k}f_{j}$ from Theorem 5.1, ∎
@@
Let $G$ be the graph as in Example 5.3. We consider the reverse lexicographic order $<$ on $R[G]$ such that
@@ 7. Algorithms
0:  The set of all stable sets of $G$, and a monomial order $<$ on $R[G]$.
0:  The reduced Gröbner basis of $K_{G}$ with respect to $<$.
1:  :=${\mathcal{F}}:=\{x_{S\setminus\{i\}}x_{\{i\}}-x_{S}x_{\emptyset}\mid i\in S\in S(G),\ |S|\geq 2\}$.
@@
3:  $\mathcal{G}_{2}:=\{g/x_{\emptyset}^{k}\mid g\in\mathcal{G},k\in{\mathbb{Z}}_{\geq 0}$, $x_{\emptyset}^{k}$ divides $g$, $x_{\emptyset}^{k+1}$ does not divide $g\}$.
4:  $\mathcal{G}_{3}:=\{x_{S_{1}}x_{S_{2}}-x_{S_{3}}x_{S_{4}}\in{\mathcal{G}}_{2}\mid S_{1},S_{2},S_{3},S_{4}\in S(G),\ S_{1}\cap S_{2}=S_{3}\cap S_{4}=\emptyset\}$.
5:  Compute the reduced Gröbner basis ${\mathcal{G}}$ of $K_{G}$ with respect to $<$ from a system of generators ${\mathcal{G}}_{3}\cup\{x_{S}x_{T}\mid S,T\in S(G),S\cap T\neq\emptyset\}$.
@@
4:     if ${\mathbf{x}}_{f_{i}}=x_{S_{1}}\cdots x_{S_{k}}$ satisfies $[d]=S_{1}\cup\cdots\cup S_{k}$ then
@@
Work with the same notation as in Theorem 6.4, that is, $\mathcal{G}=(\mathcal{G}_{1}\setminus M_{G})\cup\mathcal{G}_{2}$ is the reduced Gröbner basis of $K_{G}$ with respect to $<$. Let $\mathcal{G}_{1}\setminus M_{G}=\{{\mathbf{x}}_{p_{1}}-{\mathbf{x}}_{q_{1}},\dots,{\mathbf{x}}_{p_{t}}-{\mathbf{x}}_{q_{t}}\}$. From Theorem 1.1, $p_{j}$ and $q_{j}$ are Kempe equivalent for each $j$. Then
@@
is called a Kempe basis of $G$ with respect to $<$ if $f_{1}^{(j)}=p_{j}$, $f_{s_{j}}^{(j)}=q_{j}$, and $f_{i}^{(j)}$ is obtained from $f_{i-1}^{(j)}$ by a Kempe switching.
@@
Procedure 2. Let $f^{\prime}$ and $g^{\prime}$ be $k$-colorings of an induced subgraph $G_{1}$ of $G$. Suppose that ${\mathbf{x}}_{f^{\prime}}-{\mathbf{x}}_{g^{\prime}}={\bf x}_{h}({\mathbf{x}}_{f}-{\mathbf{x}}_{g})$. Then $f$ and $g$ are the restriction of $f^{\prime}$ and $g^{\prime}$ to some induced subgraph $G_{2}$ of $G_{1}$, respectively. Hence if $f_{0},f_{1},\ldots,f_{p}$ is a sequence of $k^{\prime}$-colorings of $G_{2}$ such that $f_{0}=f$, $f_{p}=g$, and $f_{i}$ is obtained from $f_{i-1}$ by a Kempe switching, then $f_{0}^{\prime},f_{1}^{\prime},\ldots,f_{p}^{\prime}$ satisfy $f_{0}^{\prime}=f^{\prime}$, $f_{p}^{\prime}=g^{\prime}$, and $f_{i}^{\prime}$ is obtained from $f_{i-1}^{\prime}$ by a Kempe switching, where $f_{i}^{\prime}$ is a $k$-coloring of $G_{1}$ obtained by combining $f_{i}$ with $h$.
Procedure 3. Let ${\mathbf{x}}_{p}-{\mathbf{x}}_{q}$ be a binomial in the reduced Gröbner basis ${\mathcal{G}}$ of $K_{G}$ with respect to a monomial order $<$. From Proposition 6.4, any binomial ${\mathbf{x}}_{p}-{\mathbf{x}}_{q}\in{\mathcal{G}}$ is obtained from the reduced Gröbner basis of $J_{G}$ with respect to $<$ which is computed from
$\{{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are $2$-colorings of an induced subgraph of $G$}\}$
by Buchberger Algorithm. By keeping track of the computation of Buchberger Algorithm, we can compute an expression
@@
0:  A monomial order $<$ on $R[G]$.
0:  A Kempe basis of $G$ with respect to $<$.
@@
=${\mathcal{F}}=\{{\mathbf{x}}_{f_{1}}-{\mathbf{x}}_{g_{1}},\dots,{\mathbf{x}}_{f_{t}}-{\mathbf{x}}_{g_{t}}\}\cup\{x_{S}x_{T}\mid S,T\in S(G),S\cap T\neq\emptyset\}$
@@
5:  Compute the reduced Gröbner basis ${\mathcal{G}}$ of $K_{G}$ with respect to $<$ from ${\mathcal{F}}$.
6:  $A:=\{\}$.
@@
0:  “a sequence of Kempe switchings from $f$ to $g$” or “$f\not\sim_{k}g$”.
@@
5:     By Procedure 2, from $\mathcal{K}$, find a sequence of colorings $g=g_{1},\dots,g_{t},g^{\prime}$ ($=f^{\prime}$), where $g_{i}$ is obtained from $g_{i-1}$ by a Kempe switching.

### v2

@@ 5. Examining Kempe equivalence
[…]
@@ 6. Computing Kempe classes via Kempe ideals
The set of all monomials in $M_{G}$ consists of all monomials associated with $k$-colorings of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{d}$ such that $G_{{\mathbf{a}}}$ is not an induced subgraph of $G$.
@@
Let $G$ be a graph on $[d]$, and let $\mathcal{G}_{1}$ be the reduced Gröbner basis of $J_{G}$ with respect to a monomial order << and set $\mathcal{G}_{2}=\{x_{S}x_{T}\mid S,T\in S(G),\ S\cap T\neq\emptyset\}$. Then
@@
is the reduced Gröbner basis of $K_{G}$ with respect to <<.
@@
Since $\mathcal{G}_{1}$ is a Gröbner basis of $J_{G}$, we have $J_{G}=\langle\mathcal{G}_{1}\rangle$. It follows from $M_{G}=\langle\mathcal{G}_{2}\rangle$ that $\mathcal{G}$} is a set of generators of $K_{G}=J_{G}+M_{G}$. We apply Buchberger’s Criterion to $\mathcal{G}$}. If $p,q\in\mathcal{G}_{1}\setminus M_{G}$, then the remainder of $S(p,q)$ on division by $\mathcal{G}$} is 0 since $\mathcal{G}_{1}$ is a Gröbner basis of $J_{G}$. If $p,q\in\mathcal{G}_{2}$, then both $p$ and $q$ are monomials, and hence $S(p,q)=0$. Suppose that $p\in\mathcal{G}_{2}$ and $q\in\mathcal{G}_{1}\setminus M_{G}$. Let $p=x_{S}x_{T}$ with $S,T\in S(G)$ and $S\cap T\neq\emptyset$, and let $q={\mathbf{x}}_{f}-{\mathbf{x}}_{g}$ where $f$ and $g$ are $k$-colorings of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in{\mathbb{Z}}^{d}_{\geq 0}$ and ${\rm in}_{<}(q)={\mathbf{x}}_{f}$.
@@
Case 2. (${\mathbf{x}}_{f}$ is divided by $x_{S}$.) From Lemma 6.3, ${\mathbf{x}}_{f}$ is not divided by $x_{T}$ since $q\notin M_{G}$. Then $S(p,q)=x_{T}{\mathbf{x}}_{g}$. Let $i\in S\cap T$. Since ${\mathbf{x}}_{f}$ is divided by $x_{S}$, $\pi({\mathbf{x}}_{f})=\pi({\mathbf{x}}_{g})$ is divided by $t_{i}$. Thus $\pi(x_{T}{\mathbf{x}}_{g})$ is divided by $t_{i}^{2}$. Hence $x_{T}{\mathbf{x}}_{g}$ is divided by a monomial $x_{S^{\prime}}x_{T}$ such that $i\in S^{\prime}\cap T$. Then the remainder of $S(p,q)=x_{T}{\mathbf{x}}_{g}$ on division by ${\mathcal{G}}_{2}$ is 0.
Therefore the remainder of any $S$-polynomial on division by $\mathcal{G}$} is 0. By Buchberger’s Criterion, $\mathcal{G}$} is a Gröbner basis of $K_{G}$. Since ${\mathcal{G}}_{1}$ is reduced, it is easy to see that $\mathcal{G}$} is reduced. ∎
@@
Suppose that ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in K_{G}$. We show that ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in J_{G}$. Let $\mathcal{G}$ be the reduced Gröbner basis of $K_{G}$ with respect to a monomial order <<. From Proposition 6.4, $\mathcal{G}=(\mathcal{G}_{1}\setminus M_{G})\cup\mathcal{G}_{2}$ where $\mathcal{G}_{1}$ is the reduced Gröbner basis of $J_{G}$ with respect to << and $\mathcal{G}_{2}=\{x_{S}x_{T}\mid S,T\in S(G),\ S\cap T\neq\emptyset\}$. Let ${\mathbf{x}}_{f^{\prime}}$ (resp. ${\mathbf{x}}_{g^{\prime}}$) denote the remainder of ${\mathbf{x}}_{f}$ (resp. ${\mathbf{x}}_{g}$) on division by $\mathcal{G}_{1}\setminus M_{G}$. Suppose that ${\mathbf{x}}_{f^{\prime}}\neq{\mathbf{x}}_{g^{\prime}}$. Since $f$ and $g$ are $k$-colorings of $G_{0}$, we have ${\mathbf{x}}_{f},{\mathbf{x}}_{g}\notin M_{G}$ from Lemma 6.1. From Lemma 6.3, ${\mathbf{x}}_{f^{\prime}},{\mathbf{x}}_{g^{\prime}}\notin M_{G}$ since both ${\mathbf{x}}_{f}-{\mathbf{x}}_{f^{\prime}}$ and ${\mathbf{x}}_{g}-{\mathbf{x}}_{g^{\prime}}$ belong to $J_{G}$. Thus ${\mathbf{x}}_{f^{\prime}}-{\mathbf{x}}_{g^{\prime}}$ ($\neq 0$) is the normal form of ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in K_{G}$ with respect to $\mathcal{G}$. This contradicts that $\mathcal{G}$ is a Gröbner basis of $K_{G}$. Thus ${\mathbf{x}}_{f^{\prime}}={\mathbf{x}}_{g^{\prime}}$. Then the normal form of ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}$ with respect to $\mathcal{G}_{1}\setminus M_{G}$ is zero, and hence ${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in J_{G}$. ∎
@@
Let $G$ be a graph on $[d]$ and << a monomial order on $R[G]$, and let $\{{\mathbf{x}}_{f_{1}},\dots,{\mathbf{x}}_{f_{s}}\}$ be the set of all standard monomials of degree $k$ with respect to ${\rm in}_{<}(K_{G})$. Then each $f_{i}$ is a $k$-coloring of an induced subgraph of $G$. In addition, given an induced subgraph $G^{\prime}$ of $G$,
@@
Let $\mathcal{G}$ be the reduced Gröbner basis of $K_{G}$ with respect to a monomial order <<. From Proposition 6.4, $\mathcal{G}=(\mathcal{G}_{1}\setminus M_{G})\cup\mathcal{G}_{2}$ where $\mathcal{G}_{1}$ is the reduced Gröbner basis of $J_{G}$ with respect to << and $\mathcal{G}_{2}=\{x_{S}x_{T}\mid S,T\in S(G),\ S\cap T\neq\emptyset\}$. Since each ${\mathbf{x}}_{f_{i}}$ is not divided by any monomial in $\mathcal{G}_{2}$, from Lemma 6.1, each $f_{i}$ is a $k$-coloring of an induced subgraph of $G$.
@@
Let $f$ be a $k$-coloring of $G^{\prime}$, and let ${\mathbf{x}}_{f^{\prime}}$ be the remainder of ${\mathbf{x}}_{f}$ on division by $\mathcal{G}_{1}\setminus M_{G}$. Then ${\mathbf{x}}_{f}-{\mathbf{x}}_{f^{\prime}}$ belongs to $J_{G}$. Since ${\mathbf{x}}_{f}\notin M_{G}$, we have ${\mathbf{x}}_{f^{\prime}}\notin M_{G}$ from Lemma 6.3. Thus ${\mathbf{x}}_{f^{\prime}}$ is the normal form of ${\mathbf{x}}_{f}$ with respect to $\mathcal{G}$, and hence ${\mathbf{x}}_{f^{\prime}}={\mathbf{x}}_{f_{j}}$ for some $j$. Since ${\mathbf{x}}_{f}-{\mathbf{x}}_{f_{j}}$ belongs to $J_{G}$, we have $f\sim_{k}f_{j}$ from Theorem 5.1. ∎
@@
Let $G$ be the graph as in Example 5.3. We consider the reverse lexicographic order << on $R[G]$ such that
@@ 7. Algorithms
0: The set of all stable sets of $G$, and a monomial order << on $R[G]$.
0: The reduced Gröbner basis of $K_{G}$ with respect to <<.
[…]
@@
[…]
5: Compute the reduced Gröbner basis $\mathcal{G}$} of $K_{G}$ with respect to << from a system of generators ${\mathcal{G}}_{3}\cup\{x_{S}x_{T}\mid S,T\in S(G),S\cap T\neq\emptyset\}$.
@@
4:  if ${\mathbf{x}}_{f_{i}}=x_{S_{1}}\cdots x_{S_{k}}$ satisfies $d]=S_{1}\cup\cdots\cup S_{k}$ then
@@
Work with the same notation as in Theorem 6.4, that is, $\mathcal{G}=(\mathcal{G}_{1}\setminus M_{G})\cup\mathcal{G}_{2}$ is the reduced Gröbner basis of $K_{G}$ with respect to <<. Let $\mathcal{G}_{1}\setminus M_{G}=\{{\mathbf{x}}_{p_{1}}-{\mathbf{x}}_{q_{1}},\dots,{\mathbf{x}}_{p_{t}}-{\mathbf{x}}_{q_{t}}\}$. From Theorem 1.1, $p_{j}$ and $q_{j}$ are Kempe equivalent for each $j$. Then
@@
is called a Kempe basis of $G$ with respect to << if $f_{1}^{(j)}=p_{j}$, $f_{s_{j}}^{(j)}=q_{j}$, and $f_{i}^{(j)}$ is obtained from $f_{i-1}^{(j)}$ by a Kempe switching.
@@
Procedure 2. Let $f^{\prime}$ and $g^{\prime}$ be $k$-colorings of an induced subgraph $G_{1}$ of $G$. Suppose that ${\mathbf{x}}_{f^{\prime}}-{\mathbf{x}}_{g^{\prime}}={\bf x}_{h}({\mathbf{x}}_{f}-{\mathbf{x}}_{g})$. Then $f$ and $g$ are the restrictions of $f^{\prime}$ and $g^{\prime}$ to some induced subgraph $G_{2}$ of $G_{1}$, respectively. Hence if $f_{0},f_{1},\ldots,f_{p}$ is a sequence of $k^{\prime}$-colorings of $G_{2}$ such that $f_{0}=f$, $f_{p}=g$, and $f_{i}$ is obtained from $f_{i-1}$ by a Kempe switching, then $f_{0}^{\prime},f_{1}^{\prime},\ldots,f_{p}^{\prime}$ satisfy $f_{0}^{\prime}=f^{\prime}$, $f_{p}^{\prime}=g^{\prime}$, and $f_{i}^{\prime}$ is obtained from $f_{i-1}^{\prime}$ by a Kempe switching, where $f_{i}^{\prime}$ is a $k$-coloring of $G_{1}$ obtained by combining $f_{i}$ with $h$.
Procedure 3. Let ${\mathbf{x}}_{p}-{\mathbf{x}}_{q}$ be a binomial in the reduced Gröbner basis $\mathcal{G}$} of $K_{G}$ with respect to a monomial order <<. From Proposition 6.4, any binomial ${\mathbf{x}}_{p}-{\mathbf{x}}_{q}\in{\mathcal{G}}$ is obtained from the reduced Gröbner basis of $J_{G}$ with respect to << which is computed from
[…]
by Buchberger’s Algorithm. By keeping track of the computation of Buchberger’s Algorithm, we can compute an expression
@@
0: A monomial order << on $R[G]$.
0: A Kempe basis of $G$ with respect to <<.
@@
${\mathcal{F}}=\{{\mathbf{x}}_{f_{1}}-{\mathbf{x}}_{g_{1}},\dots,{\mathbf{x}}_{f_{t}}-{\mathbf{x}}_{g_{t}}\}\cup\{x_{S}x_{T}\mid S,T\in S(G),S\cap T\neq\emptyset\}$
@@
5: Compute the reduced Gröbner basis $\mathcal{G}$} of $K_{G}$ with respect to << from $\mathcal{F}$}.
[…]
@@
Although Algorithm 7.1 requires the enumeration of all stable sets and the computation of a Gröbner basis, and thus is not computationally feasible for large graphs, its merit lies in providing a uniform algebraic framework. In particular, the method is mechanically applicable and conceptually clarifies the structure of the graph from an abstract algebraic viewpoint.
@@
0: “a sequence of Kempe switchings from $f$ to gg” or $f\not\sim_{k}g$”.
@@
5:  By Procedure 2, from $\mathcal{K}$, find a sequence of colorings $g=g_{1},\dots,g_{t},g^{\prime}$ ($f^{\prime}$), where $g_{i}$ is obtained from $g_{i-1}$ by a Kempe switching.